# MFC 5.0: An exascale many-physics flow solver

### v1

@@ 4.2.1 General characteristic boundary conditions
4.2.2 WENO-Z, TENO, and seventh-order non-uniform reconstruction
The WENO-Z scheme [64] enhances classical WENO methods by improving their spectral properties while being faster than WENO-M. The $(2k-1)$-th order WENO-Z weight $\omega_{r}^{(\mathrm{Z})}$ for the $r$-th stencil is calculated as
@@ 4.2.2 WENO-Z, TENO, and seventh-order non-uniform reconstruction
where $d_{r}$ are the ideal weights, and $\tau$ is the global smoothness indicator. The parameter $q$ influences the convergence rate at the critical points and affects spectral properties. This formulation is valid for schemes of order 3, 5, and 7.
The Targeted Essentially Non-Oscillatory (TENO) scheme further reduces numerical dissipation by implementing a stencil selection process akin to the ENO method [65]. A $K$-th order accurate TENO scheme uses $K-2$ candidate stencils of increasing width. The TENO equations are
@@
where $C_{\mathrm{T}}$ is a smoothness threshold determining stencil activation and $\tau$ follows a similar definition as that of WENO-Z.
For seventh-order reconstruction on non-uniform grids, the explicit equations are omitted due to their length. Instead, we detail the process below. The reconstruction polynomial of each stencil using cell-average values is obtained through the Lagrange interpolation method [66]:
$p(x)=\sum_{j=0}^{k-1}\hskip 5.0pt\underbrace{\Delta x_{j}\sum_{m=j+1}^{k}\left(\frac{\sum_{\begin{subarray}{c}l=0\\ l\neq m\end{subarray}}^{k}\prod_{\begin{subarray}{c}q=0\\ q\neq m,\,l\end{subarray}}^{k}\left(x-x_{q-\frac{1}{2}}\right)}{\prod_{\begin{subarray}{c}l=0\\ l\neq m\end{subarray}}^{k}\left(x_{m-\frac{1}{2}}-x_{l-\frac{1}{2}}\right)}\right)}_{\mbox{$c_{j}(x)$}}\hskip 5.0ptf_{j},$
where $\Delta x_{j}:=x_{j+1/2}-x_{j-1/2}$. We compute $p_{r}:=p(k-r-1/2)$ for the right flux of the $r$-th stencil. To reduce computational cost, we rewrite $p(x)$ in terms of $\Delta f_{j}:=f_{j+1}-f_{j}$, so $p(x)=\sum_{j=0}^{k-2}\hat{c}_{j}(x)\Delta f_{j}+f_{0}$. The coefficients $\hat{c}_{r,j}:=\hat{c}_{j}(k-r-1/2)$ are precomputed at the start of the program.
The smoothness indicators are used to derive $\hat{b}_{r,j,m}$ for each $\Delta f_{j}\Delta f_{k}$ cross term
$\beta_{r}$ $=\sum_{l=1}^{k-1}\int_{x_{k-r-\frac{3}{2}}}^{x_{k-r-\frac{1}{2}}}\Delta x^{2l-1}\left(\frac{\partial^{l}p(x)}{{\partial x}^{l}}\right)^{2}\text{d}x$
$=\sum_{j=0}^{k-2}\sum_{m=j}^{k-2}\hskip 5.0pt\underbrace{\left(2-\delta_{jm}\right)\sum_{l=1}^{k-1}\int_{x_{k-r-\frac{3}{2}}}^{x_{k-r-\frac{1}{2}}}\Delta x^{2l-1}\frac{\partial^{l}\hat{c}_{j}(x)}{{\partial x}^{l}}\frac{\partial^{l}\hat{c}_{m}(x)}{{\partial x}^{l}}\text{d}x}_{\mbox{$\hat{b}_{r,j,m}$}}\hskip 5.0pt\Delta f_{j}\Delta f_{m}.$
Again, $\hat{b}_{r,j,m}$ can be precomputed. By using $\Delta f$, we reduce the number of terms from $k$ to $k-1$ for $f_{r}$, and from $k(k+1)/2$ to $(k-1)k/2$ for $\beta_{r}$, decreasing the computational operations required.
@@ 4.3.1 GPU Offloading with OpenACC on NVIDIA and AMD Hardware
Vendor portable GPU offloading on AMD and NVIDIA hardware is implemented using OpenACC [72]. OpenACC is a directive-based tool that allows the developer to generate GPU kernels by adding directive clauses around developer-identified regions of parallelism in the code. OpenACC is supported by NVIDIA’s NVHPC compiler for use with NVIDIA GPUs and HPE’s Cray Compiler Environment for use with both NVIDIA and AMD GPUs. Other compilers, like GNU 14+ and Flang 18+, support OpenACC, though their relative immaturity at time of writing makes them insufficient to support MFC.
LABEL:lst:kernel shows an example of a typical GPU kernel in MFC. The outer loops over j, k, and l iterate through $\mathcal{O}(100)$ elements each, corresponding to the grid cells in the MPI subdomain. The inner loop over i iterates through each equation in the current model. For two-phase flow problems this loop spans $\mathcal{O}(1)$ elements. When additional features like hypo- and hyperelasticity or chemistry are added to the problem, the extent of this inner loop can be $\mathcal{O}(10)$.
@@ 4.3.2 Performance on various CPU, GPU, and APU architectures
MFC 5.0 has been benchmarked on a broad range of CPU, GPU, and APU (superchip) devices. LABEL:t:serialperf shows report the single-device performance in terms of its grind time, which is computed as time per grid point, PDE, and right-hand side evaluation. The grind times are calculated for one GPU or APU device or one CPU socket. The benchmarks were performed under a similar strategy as that used on published testbed reports [78]. These quantities are for a typical, representative compressible multi-component problem: 3D, inviscid, 5-equation model problem with two advected species (8 PDEs) and 8M grid points (158-cubed 3D uniform grid). The equation is solved via fifth-order accurate WENO finite volume reconstruction and the HLLC approximate Riemann solver. This case is in the MFC source code under examples/3D_performance_test. One can execute it via
@@
One can execute this command for CPU cases, which builds an optimized version of the code for this case, then executes it. For benchmarking GPU devices, one will likely want to use -n <num_gpus> where the usual case for single GPU device benchmarking leaves <num_gpus> as unity. Similar performance is also seen for other problem configurations, such as the Euler equations (4 PDEs). All results are for the compiler that gave the best performance, including AOCC, Intel, GCC, CCE, and NVHPC.
@@
Roofline performance is presented for the most expensive kernels on GPU offerings from NVIDIA and AMD in fig. 4. At the time of writing, roofline analysis is unavailable on the MI300A APU, so it remains absent. NVHPC 24+ compilers and NVIDIA Nsight Compute profilers gather roofline results on NVIDIA hardware and CCE 18+ with Omniperf gathers roofline results on the AMD MI250X GPU. The WENO and approximate Riemann kernels achieve 37% and 14% of peak performance on the NVIDIA A100. On the NVIDIA GH200, the WENO and approximate Riemann kernels achieve 24% and 5% of peak compute performance. The AMD MI250X achieves 22% of peak compute performance in the WENO kernel and 3% of peak performance in the approximate Riemann problem.
Figure 4: Roofline performance of the fifth-order accurate WENO reconstruction and HLLC approximate Riemann solve kernels on the AMD MI250X and NVIDIA GH200 and A100.
The breakdown of time spent in the two most expensive kernels, time spent packing arrays, and time spent doing all other computations and communication is shown in fig. 5. The WENO and approximate Riemann kernels take up just over 50% of the total time on NVIDIA GPUs. On AMD GPU devices, the time spent in the WENO and approximate Riemann kernels is similar to NVIDIA devices. Still, these kernels make up about a third of the total wall time due to an increase in time spent packing arrays for coalesced memory access. Profiles show a high number of cache misses on the AMD MI250X, which is suspected to be due to its smaller caches when compared to its NVIDIA counterparts.
@@ 4.3.3 Exascale capabilities and beyond
MFC efficiently scales to tens of thousands of GPUs on the world’s fastest computers. Figure 6 shows the weak scaling performance of MFC on OLCF Summit with NVIDIA V100 GPUs, OLCF Frontier with AMD MI250X GPUs, and LLNL El Capitan with AMD MI300A APUs. MFC scales from 128 to 13824 NVIDIA V100 GPUs on OLCF Summit with 97% efficiency and from 128 to 65536 AMD MI250X GCDs (64 to 32768 MI250X GPUs) on OLCF Frontier with 95% efficiency.
@@
The simple communication patterns in MFC also result in high strong scaling efficiencies. Figure 7 shows the strong scaling performance of MFC on OLCF Summit and OLCF Frontier and compares weak scaling with and without GPU-aware MPI on both systems. The problem sizes in fig. 7 (a) correspond to approximately 100% and 50% usage of available GPU memory per device in the base case. For the case with 8M grid cells per device in the base case, strong scaling efficiencies of 79% and 51% are observed on OLCF Summit when increasing the device count by a factor of eight and 16 without GPU-aware MPI. For the case with 32 million grid cells per device in the base case, strong scaling efficiencies of 81% and 77% when increasing the device count by a factor of 8 and 16 without GPU-aware MPI. The higher efficiencies seen with the AMD MI250X on OLCF Frontier are partly due to the increased computation-to-communication ratio of the larger problem size but also the improved internode communication on Frontier. When GPU-aware MPI is enabled, the strong scaling efficiencies on Summit increase to 84% and 60% when increasing the device count by a factor of 8 and 16. On Frontier, the efficiencies increase to 96% and 92% when device count is increased by a factor of 8 and 16.

### v3

@@ 4.2.1 General characteristic boundary conditions
4.2.2 WENO-Z, TENO, and high-order non-uniform reconstruction
The WENO-Z scheme [68] enhances classical WENO methods by improving their spectral properties while being faster than WENO-M. The $(2k-1)$-th order WENO-Z weight $\omega_{r}^{(\mathrm{Z})}$ for the $r$-th stencil is computed as
@@ 4.2.2 WENO-Z, TENO, and high-order non-uniform reconstruction
where $d_{r}$ are the ideal weights, $\tau$ is the global smoothness indicator, and the parameter $q$ (typically set to $1$ for fifth-order reconstruction) influences the convergence rate at the critical points and the spectral properties. This formulation applies to schemes of order 3, 5, and 7.
The Targeted Essentially Non-Oscillatory (TENO) scheme further reduces numerical dissipation by employing a stencil selection process similar to the ENO method [69]. A $K$-th order accurate TENO scheme uses $K-2$ candidate stencils of increasing width. Its equations are given by
@@
where $C_{T}$ is a smoothness threshold for stencil activation and $\tau$ is defined as in the WENO-Z scheme.
The explicit expressions for the seventh-order WENO reconstruction in our code are omitted here due to their considerable length. Instead, we present a generalized approach for reconstructing a $(2k-1)$-order scheme from cell averages on non-uniform grids. For each stencil, the reconstruction polynomial is obtained using the Lagrange interpolation method [70]:
$p(x)=\sum_{j=0}^{k-1}\hskip 5.0pt\underbrace{\Delta x_{j}\sum_{m=j+1}^{k}\left(\frac{\sum_{\begin{subarray}{c}l=0\\ l\neq m\end{subarray}}^{k}\prod_{\begin{subarray}{c}q=0\\ q\neq m,\,l\end{subarray}}^{k}\left(x-x_{q-\frac{1}{2}}\right)}{\prod_{\begin{subarray}{c}l=0\\ l\neq m\end{subarray}}^{k}\left(x_{m-\frac{1}{2}}-x_{l-\frac{1}{2}}\right)}\right)}_{\mbox{$c_{j}(x)$}}\hskip 5.0ptf_{j}.$
Here, $x_{j}$ denotes cell-center location, $x_{j+1/2}$ the cell-boundary location, and $\Delta x_{j}:=x_{j+1/2}-x_{j-1/2}$ is the cell width. For each stencil $r=0,1,\dots,k-1$, the $k$-cell stencil spans the interval from $x_{-1/2}$ to $x_{k-1/2}$. Consequently, the overall $(2k-1)$-cell stencil is centered at $x_{k-r-1}$, and the right flux is $p_{r}:=p(x_{k-r-1/2})$.
To reduce computational cost, we rewrite $p(x)$ in terms of the differences $\Delta f_{j}:=f_{j+1}-f_{j}$, so that $p(x)=\sum_{j=0}^{k-2}\hat{c}_{j}(x)\Delta f_{j}+f_{0}$. For flux evaluation at $x_{k-r-1/2}$, the coefficients $\hat{c}_{r,j}:=\hat{c}_{j}(x_{k-r-1/2})$ for each $\Delta f_{j}$ and stencil $r$ are precomputed at the start of the program.
The smoothness indicator for stencil $r$ is defined by
$\beta_{r}=\sum_{l=1}^{k-1}\Delta x_{k-r-1}^{2l-1}\int_{x_{k-r-\frac{3}{2}}}^{x_{k-r-\frac{1}{2}}}\left(\frac{\partial^{l}p(x)}{{\partial x}^{l}}\right)^{2}\text{d}x.$
Rewriting this in terms of the difference formulation yields
$\beta_{r}=\sum_{j=0}^{k-2}\sum_{m=j}^{k-2}\hskip 5.0pt\underbrace{\left(2-\delta_{jm}\right)\sum_{l=1}^{k-1}\Delta x_{k-r-1}^{2l-1}\int_{x_{k-r-\frac{3}{2}}}^{x_{k-r-\frac{1}{2}}}\frac{\partial^{l}\hat{c}_{j}(x)}{{\partial x}^{l}}\frac{\partial^{l}\hat{c}_{m}(x)}{{\partial x}^{l}}\text{d}x}_{\mbox{$\hat{b}_{r,j,m}$}}\hskip 5.0pt\Delta f_{j}\Delta f_{m},$
where $\delta_{jm}$ is the Kronecker delta. The coefficients $\hat{b}_{r,j,m}$ for each cross term $\Delta f_{j}\Delta f_{m}$ and stencil $r$ are precomputed. Reformulating in terms of $\Delta f$ reduces the number of terms from $k$ to $k-1$ for $p_{r}$, and from $k(k+1)/2$ to $(k-1)k/2$ for $\beta_{r}$, thereby reducing the computational cost.
@@ 4.2.4 Low-Mach number treatment
4.2.5 HLLD Riemann Solver
For the MHD equations, the HLLD (Harten–Lax–van Leer–Discontinuities) Riemann solver [76] build on the HLL framework to resolve five of the seven characteristic waves, thereby reducing numerical dissipation and improving accuracy.
In this approach, the Riemann fan is divided into four intermediate states $\mathbf{U}_{\mathrm{L}}^{*}$, $\mathbf{U}_{\mathrm{L}}^{**}$, $\mathbf{U}_{\mathrm{R}}^{**}$, and $\mathbf{U}_{\mathrm{R}}^{*}$. These states are separated by two Alfvén waves ($S_{\mathrm{L}}^{*}$ and $S_{\mathrm{R}}^{*}$) and a middle contact discontinuity ($S_{\mathrm{M}}$). The fastest left and right waves ($S_{\mathrm{L}}$ and $S_{\mathrm{R}}$) form the outer boundaries of the entire Riemann fan.
Similar to the HLL solver, $S_{\mathrm{L}}$ and $S_{\mathrm{R}}$ are approximated by
$S_{\mathrm{L}}=\min(u_{\mathrm{L}}-c_{f,\mathrm{L}},u_{\mathrm{R}}-c_{f,\mathrm{R}})\quad\text{and}\quad S_{\mathrm{R}}=\max(u_{\mathrm{L}}+c_{f,\mathrm{L}},u_{\mathrm{R}}+c_{f,\mathrm{R}}).$
The HLLD solver enforces two key constraints: normal velocity and total pressure remain constant across the Riemann fan:
$u_{\mathrm{L}}^{*}=u_{\mathrm{L}}^{**}=u_{\mathrm{R}}^{**}=u_{\mathrm{R}}^{*}=S_{\mathrm{M}},$
$p_{\mathrm{T},\mathrm{L}}^{*}=p_{\mathrm{T},\mathrm{L}}^{**}=p_{\mathrm{T},\mathrm{R}}^{**}=p_{\mathrm{T},\mathrm{R}}^{*}=p^{*}_{\mathrm{T}}.$
The middle wave speed is calculated as:
$S_{\mathrm{M}}=\frac{(S_{\mathrm{R}}-u_{\mathrm{R}})\rho_{\mathrm{R}}u_{\mathrm{R}}-(S_{\mathrm{L}}-u_{\mathrm{L}})\rho_{\mathrm{L}}u_{\mathrm{L}}-p_{\mathrm{T},\mathrm{R}}+p_{\mathrm{T},\mathrm{L}}}{(S_{\mathrm{R}}-u_{\mathrm{R}})\rho_{\mathrm{R}}-(S_{\mathrm{L}}-u_{\mathrm{L}})\rho_{\mathrm{L}}}.$
For the outer intermediate states, we first compute the density:
$\rho_{\alpha}^{*}=\rho_{\alpha}\frac{S_{\alpha}-u_{\alpha}}{S_{\alpha}-S_{\mathrm{M}}},$
where $\alpha=\mathrm{L}$ or $\mathrm{R}$. For brevity, explicit expressions for tangential velocities ($v_{\alpha}^{*}$, $w_{\alpha}^{*}$) and tangential magnetic fields ($B_{y,\alpha}^{*}$, $B_{z,\alpha}^{*}$), as well as the energy computation, are not shown here but can be found in the original paper [76].
The Alfvén wave speeds are determined by:
$S_{\mathrm{L}}^{*}=S_{\mathrm{M}}-\frac{|B_{x}|}{\sqrt{\rho_{\mathrm{L}}^{*}}},\quad S_{\mathrm{R}}^{*}=S_{\mathrm{M}}+\frac{|B_{x}|}{\sqrt{\rho_{\mathrm{R}}^{*}}}.$
For inner states, the tangential velocity is:
$v^{**}=\frac{\sqrt{\rho_{\mathrm{L}}^{*}}v_{\mathrm{L}}^{*}+\sqrt{\rho_{\mathrm{R}}^{*}}v_{\mathrm{R}}^{*}+(B_{y,\mathrm{R}}^{*}-B_{y,\mathrm{L}}^{*})\,\text{sign}(B_{x})}{\sqrt{\rho_{\mathrm{L}}^{*}}+\sqrt{\rho_{\mathrm{R}}^{*}}}.$
Similarly, the corresponding expressions for $w^{**}$, $B_{y}^{**}$, and $B_{z}^{**}$ in the inner states, along with the associated energy calculation, follow a similar pattern and can be found in [76].
The numerical flux is chosen based on the region of the Riemann fan that contains the cell interface:
$\mathbf{F}_{\mathrm{HLLD}}=\begin{cases}\mathbf{F}_{\mathrm{L}}&\text{if }S_{\mathrm{L}}>0,\\ \mathbf{F}_{\mathrm{L}}^{*}&\text{if }S_{\mathrm{L}}\leq 0\leq S_{\mathrm{L}}^{*},\\ \mathbf{F}_{\mathrm{L}}^{**}&\text{if }S_{\mathrm{L}}^{*}\leq 0\leq S_{\mathrm{M}},\\ \mathbf{F}_{\mathrm{R}}^{**}&\text{if }S_{\mathrm{M}}\leq 0\leq S_{\mathrm{R}}^{*},\\ \mathbf{F}_{\mathrm{R}}^{*}&\text{if }S_{\mathrm{R}}^{*}\leq 0\leq S_{\mathrm{R}},\\ \mathbf{F}_{\mathrm{R}}&\text{if }S_{\mathrm{R}}<0.\end{cases}$
@@ 4.3.1 GPU Offloading with OpenACC on NVIDIA and AMD Hardware
Vendor portable GPU offloading on AMD and NVIDIA hardware is implemented using OpenACC [77]. OpenACC is a directive-based tool that allows the developer to generate GPU kernels by adding directive clauses around developer-identified regions of parallelism in the code. OpenACC is supported by NVIDIA’s NVHPC compiler for use with NVIDIA GPUs and HPE’s Cray Compiler Environment for use with both NVIDIA and AMD GPUs. Compilers like GNU 14 and Flang 18 support OpenACC, though their relative immaturity at time of writing makes them insufficient to support MFC.
LABEL:lst:kernel shows an example of a typical GPU kernel in MFC. The outer loops over j, k, and l iterate through $\mathcal{O}(100)$ elements each, corresponding to the grid cells in the MPI subdomain. The inner loop over i iterates through each equation in the current model. For two-phase flow problems, this loop spans $\mathcal{O}(1)$ elements. When additional features like hypo- and hyperelasticity or chemistry are added to the problem, the extent of this inner loop can be $\mathcal{O}(10)$.
@@ 4.3.2 Performance on various CPU, GPU, and APU architectures
MFC 5.0 has been benchmarked on a broad range of CPU, GPU, and APU (superchip) devices. LABEL:t:serialperf reports the single-device performance in terms of its grind time, which is computed as time per grid point, PDE, and right-hand side evaluation. The grind times are calculated for one GPU or APU device or one CPU socket. The benchmarks were performed under a similar strategy as that used on published testbed reports [83]. These quantities are for a typical, representative compressible multi-component problem: 3D, inviscid, 5-equation model problem with two advected species (8 PDEs) and 8M grid points (158-cubed 3D uniform grid). The equation is solved via fifth-order accurate WENO finite volume reconstruction and the HLLC approximate Riemann solver. This case is in the MFC source code under examples/3D_performance_test. One can execute it via
@@
This command can be executed for CPU cases. It builds an optimized version of the code for this case and then executes it. For benchmarking GPU devices, one will likely want to use -n <num_gpus> where the usual case for single GPU device benchmarking leaves <num_gpus> as unity. Similar performance is also seen for other problem configurations, such as the Euler equations (4 PDEs). All results are for the compiler that gave the best performance, including AOCC, Intel, GCC, CCE, and NVHPC.
@@
Figure 4 shows the roofline performance for the most expensive kernels on NVIDIA and AMD accelerators. At the time of writing, roofline analysis is unavailable on the MI300A APU, and so is not shown. NVHPC 24.5 compilers and NVIDIA Nsight Compute profilers produce roofline results on NVIDIA devices, and CCE 18 with Omniperf profile the roofline on the AMD MI250X GPU. The WENO and approximate Riemann kernels achieve 37% and 14% of peak performance on the NVIDIA A100. On the NVIDIA GH200, the WENO and approximate Riemann kernels achieve 24% and 5% of peak compute performance. The AMD MI250X achieves 22% of peak compute performance in the WENO kernel and 3% of peak performance in the approximate Riemann problem.
Figure 4: The roofline performance of the fifth-order accurate WENO reconstruction and HLLC approximate Riemann solve kernels on the AMD MI250X, NVIDIA GH200, and NVIDIA A100.
The breakdown of time spent in the two most expensive kernels, time spent packing arrays, and time spent doing all other computations and communication is shown in fig. 5. The WENO and approximate Riemann kernels take up just over 50% of the total time on NVIDIA GPUs. On AMD GPU devices, the time spent in the WENO and approximate Riemann kernels is similar to that of NVIDIA devices. Still, these kernels account for about a third of the total wall time due to an increase in the time spent packing arrays for coalesced memory access. The profiles show a high number of cache misses on the AMD MI250X, which we expect is due to the smaller device caches than those of the shown NVIDIA device.
@@ 4.3.3 Exascale capabilities and beyond
MFC efficiently scales to tens of thousands of GPUs on the world’s fastest computers. Figure 6 shows the weak scaling performance of MFC on OLCF Summit (now retired) with NVIDIA V100 GPUs, OLCF Frontier with AMD MI250X GPUs, and LLNL El Capitan with AMD MI300A APUs. MFC scales to 13824 NVIDIA V100 GPUs on OLCF Summit with 97% efficiency, to 65536 AMD MI250X GCDs (32768 MI250X GPUs) on OLCF Frontier with 95% efficiency, and to 196608 AMD MI300As XCDs (32768 MI300A APUs) on LLNL El Capitan.
@@
The communication patterns in MFC also result in high strong scaling efficiencies. Figure 7 shows the strong scaling performance of MFC on OLCF Summit and OLCF Frontier and compares weak scaling with and without GPU-aware MPI. The problem sizes in fig. 7 (a) correspond to approximately 100% and 50% usage of available GPU memory per device in the base case. In the base case, with 8M grid cells per device, strong scaling efficiencies of 79% and 51% are observed on OLCF Summit when increasing the device count by a factor of 8 or 16 without GPU-aware MPI. For the case with 32 million grid cells per device in the base case, strong scaling efficiencies of 81% and 77% when increasing the device count by a factor of 8 and 16 without GPU-aware MPI. The higher efficiencies seen with the AMD MI250X on OLCF Frontier are partly due to the larger problem size’s increased computation-to-communication ratio and the improved internode communication on Frontier. With GPU-aware MPI, the strong scaling efficiencies on OLCF Summit increase to 84% and 60% when increasing the device count by a factor of 8 and 16. On Frontier, the efficiencies are 96% and 92% when the device count is increased by the same factors.